The time taken for effusion of 64 mL of oxygen will be same as the time  taken for the effusion of which  of the following gases under identical conditions:


a) 64 mL of H2     


    b) 100 mL of N2


    c) 64 mL of CO   


d)45.24 mL of SO

Grade:12

1 Answers

Gaurav
9 years ago
\frac{V_{SO_2}}{V_{O_2}}=\sqrt{\frac{M_{O_2}}{M_S_O_2}}
or
V_{SO_2}=64\sqrt{\frac{32}{64}} =45.25 mL